2 numbers is 26 and 1 number is 12 more than the other. find the number
Answer by checkley77(12844)

x+y=26
x-y=12 add
----------------
2x=38
x=38/2
x=19 ans.
19-y=12
-y=12-19
-y=-7 ans.
Proof:
19+7=26